OpenStudy (anonymous):

An equation of a circle is x^2 + y^2 + 10x – 6y + 18 = 0. Part 1: Show all your work in determining the center and radius of this circle. Part 2: In complete sentences, explain the procedure used.

OpenStudy (anonymous):

use completing the square.

myininaya (myininaya):

Put your x guys together and put your y's guys together. You need to be able to complete the square. Recall (m+n)^2=m^2+2mn+n^2 So if we have a^2+ja+k and we want to find what k is that completes the square: We divide j/2 then square that result. This will be k. k=(j/2)^2 So if we have a^2+ja+(j/2)^2 we can rewrite this as (a+j/2)^2 Or if we have m^2+2mn+? ?=(2n/2)^2=(n)^2=n^2 So m^2+2mn+n^2=(m+n)^2 like we said above.

OpenStudy (anonymous):

get the constants in one side. then complete the square for x then for y